OpenStudy (didierr):

A computer program changed one number into another. The program changes 12 into 9, 20 into 15, and 100 into 75. What number will 60 be changed into? 12 → program → 9 20 → program → 9 → 9 100 → program → 75 60 → program → ?

OpenStudy (janu16):

still need help?

OpenStudy (janu16):

now you have the equation 60/x = 4/3, and all you have to do is simply solve for x. 60/x = 4/3 (multiply by x) 60 = 4x/3 (multiply by 3) 180 = 4x (divide by 4) x = 45

OpenStudy (janu16):

@didierr